5. What is the pH of a solution prepared by mixing 25.00 mL. of 0.10 M methylamine, CH3NH2, with 25.00 mL of 0.10 M methylammonium chloride, CH3NH3C1? Assume that the volume of the solutions are additive and that Kb-3.70 x 10-4 for methylamine. A) 10.27 B) 10.57 C) 10.87 D) 11.78 6. Which statement about buffers is true? A) Buffers have a pH-7 B) Buffers consist of a strong acid and its conjugate base. C) A buffer does not change pH on addition of a strong acid or strong base. D) Buffers resist change in pH upon addition of small amounts of strong acid or strong base. 7. What is the pH of a buffered system made by dissolving 17.42 g of KH2PO4 and 20.41 g of K2HPO4 in water to give a volume of 200.0 mL? The Ka2 for dihydrogen phosphate is 6.2 × 10-8 and the equilibrium reaction of interest is H2P04-(aq) + H2O(l) = H30+(aq) + HPO4-(aq). A) 7.03 B) 7.17 C) 7.38 D) 7.58 8. TRIS ((HOCH2)3CNH2) is one of the most common buffers used in biochemistry. A solution is prepared by adding enough TRIS and 12 M HCI(ag) to give 1.00 L of solution with [TRIS]-0.30 M and ITRISH+]-0.60 M. What is the pH of this buffered system if the pKb is 5.92? A) 5.92 B) 6.22 C)7.78 D) 8.08 9. A buffer solution is prepared by dissolving 27.22 g of KH2PO4 and 3.37 g of KOH in enough water to make 0.100 L of solution. What is the pH of the H2PO4-/HPO42-buffer if the Ka2 6.2 x 10-8? A) 6.84
Answer #1

N2= Akta Lem then of CH3NH2 in mixtures? V2 final Volume of ta mix 5+25)m-50ml. (m) =0.05m 5b 2 Similayfr meleyl am methyl am

0 b64 pH=7.21-0.04 = 717 o t рон = 59114630) CTRi5T 6.60 o 30) Рон = 6, 22